6(3-z)/5=-z
3(x-5)/2=2(x+5)/3

Answer by rfer(16322)

18-6z=-5z
18=-5z+6z
18=z
-----------
3x-15/2=2x+10/3
3(3x-15)=2(2x+10)
9x-45=4x+20
9x-4x=20+45
5x=65
x=65/5
x=13